(b) How does the anewer change it the coefficient of kinetic tiction is Is p.1307 (use the aman pustin acrteofmass m . 2 Okgat a spend at v-0E75 m/s encounters er length 1-0.65 m, as the naure been inne coene ere·rinenc meter, betwsen the crate and rough surtace is a 234 ana ne exerts a constart 意 TOSHIBA
Answer #1

given

m = 92 kg

v = 0.875 m/s

l = 0.65 m

k = 0.354

F = 282 N

a. on the rough surface

net force = F - kN

N = mg

hence

Fnet = F - kmg = 282 - 0.354*92*9.81 = -37.49208 N ( opposite to the direction of motion of the cart)

b. work done on crate while it is on rough surface = W

W = F*l = 282*0.65 = 183.3 J

c. final speed = u

from conservation of energy

-37.49208*0.65 = 0.5*92(u^2 - 0.875^2)

u = 0.48 m/s